OpenStudy (anonymous):

If y^3 = 25 x^2, determine \frac{dx}{dt} when x = 5 and \frac{dy}{dt} = -2 I'm new to this topic and i'm confused since its y^3 instead of just y, thanks in advance

OpenStudy (anonymous):

Okay this is slightly harder, but the same concepts apply. \[y^3=25x^2\]Differentiating both sides yields: \[3y^2\frac{dy}{dx}=50x\]Divide both sides by 3y^2 to give us our dy/dx: \[\frac{dy}{dx}=\frac{50x}{3y^2}\] We can find our answer using this equation:\[\frac{dx}{dt} = \frac{dx}{dy} * \frac{dy}{dt} \]We know \[\frac{dx}{dy} = \frac{1}{\frac{dy}{dx}} = \frac{3y^2}{50x} \]\[\frac{dx}{dy} (5) = \frac{3y^2}{250}\]and \[\frac{dy}{dt} = -2 \] Hence: \[\frac{dx}{dt} = \frac{-6y^2}{250} \]

OpenStudy (anonymous):

@tom982 I have to reply in webwork and it doesn't accept the answer.. or is there more steps to it?

jimthompson5910 (jim_thompson5910):

if x = 5, then you can plug that into y^3 = 25x^2 and solve for y once you know y, you can plug that into \[\Large \frac{dx}{dt} = \frac{-6y^2}{250}\] don't forget to reduce or simplify fully as much as possible

OpenStudy (anonymous):

Thanks Jim. I missed a pretty crucial step!

OpenStudy (anonymous):

ohhh i see, first solve for y then plug in, thanks guys @tom982 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

np